# MiraGe: Editable 2D Images using Gaussian Splatting

### v1

@@ 1 Introduction
Figure 1: MiraGe encodes 2D images with parameterized Gaussians, enabling high-quality reconstruction and real-life-like modifications. Selected part of image can be transformed in 3D space, creating a 3D effect, with a physics engine controlling movement and interactions.
@@
Gaussian Splatting (3DGS) framework models the structure of a 3D scene using Gaussian components (Kerbl et al., 2023). In the 2D domain, GaussianImage (Zhang et al., 2024) has shown promising results in image reconstruction by efficiently encoding images in the 2D space, with a strong focus on model efficiency and reduced training time. Regrettably, the GaussianImage lacks the capability for human-driven adjustments of scene objects, which is a crucial aspect of 3DGS. In our work, we address this by introducing the MiraGe [^1] https://github.com/waczjoan/MiraGe/blob/main/README.md model, which encodes 2D images through the lens of human perception, bridging the gap between 2D image representation and 3D spatial understanding (see Fig. 2).
@@
Building on the foundational idea that humans intuitively can perform transformations on photographs–primarily through affine transformations and bending them beyond the 2D plane–we introduce a novel approach using flat Gaussians with GaMeS parameterization (Waczyńska et al., 2024b). This allows our model to facilitate image editing within 2D and 3D spaces. In particular, our framework enables changes in the image’s perspective with minimal effort. Perspective adjustments, typically complex, are simplified by direct editing in the third dimension (see Fig. 3).
Figure 3: MiraGe simplifies intuitive editing of image, allowing transformations such as adjusting the tilt of a hand with minimal complexity. This is achieved by modifying the object along the third dimension.
In addition to classical edits, our model has the unique capability of interfacing with physics engines, enabling applications that enhance the realism and immersiveness of animations (Jiang et al., 2024). MiraGe treats the physics engine as a black box and offers three distinct methods for controlling Gaussians, i.e., 2D, Amorphous and Graphite. For 2D representation (2D-MiraGe) we used Taichi_elements, for 3D representation (Amorphous-MiraGe, Graphite-MiraGe) we use Blender. This flexibility makes our model highly applicable to various fields, such as computer graphics for populating spatial interfaces, where realistic, physics-factual 2D animations can be incorporated (Tadeja et al., 2023).
@@
Since high-quality image reconstruction is critical in animation, we compared MiraGe with other models, showing our state-of-the-art performance in the image reconstruction task. Our results demonstrate that the model operates in real-time, though this comes with a trade-off in terms of model size (see Fig. 5).
Figure 4: Two images were encoded using the MiraGe model on distinct planes within a 3D space. This setup allows for seamless integration of the encoded images, resulting in a collage-like composition. Moreover, the model facilitates editing capabilities, as illustrated here with modifications to the background image (the rear plane).
It is worth highlighting that flat 3D Gaussians can be utilized for 2D images in four distinct scenarios, with modifications that emphasize how controlling the Gaussians during training affects the perspective of viewing each image (see Fig. 6).
The following constitutes a list of our key contributions:
  • We introduce the MiraGe model, which represents 2D images using flat Gaussian components, achieving state-of-the-art reconstruction quality.
@@
  • We integrate MiraGe with a physics engine, enabling physics-based modifications and interactions for both 2D and 3D environments
Figure 5: MiraGe outperforms other models in the image reconstruction task. The following columns present ground truth image, GaussianImage reconstruction and MiraGe reconstruction. The bottom row graphics show the difference between the ground truth image and the outcomes of subsequent methods.
@@ 2 Related Works
One rapidly growing area in image reconstruction is Implicit Neural Representations (INRs), which have attracted significant attention for their ability to model continuous signals, such as images, through neural networks (Klocek et al., 2019). INRs encode spatial coordinates and map them to corresponding values, such as RGB color, allowing for highly compact and efficient representations (Xie et al., 2022). This has led to the development of several specialized models for image INR, such as SIREN (Sitzmann et al., 2020a) with the novelty of sine used as a periodic activation function to tackle the problem of complex image signals. Fourier feature mapping was proposed in (Tancik et al., 2020) as another answer to the difficulty of aligning the multilayer perceptron (MLP) predictions with high-frequency pictures. Interestingly, authors of WIRE (Saragadam et al., 2023a) have leveraged continuous complex Gabor wavelets to capture visual signals with decent quality. The growing field of research resulted in further improvements of already existing solutions, e.g. in (Liu et al., 2024), certain limitations of SIREN, namely the arising capacity-convergence gap, were successfully alleviated with the idea of variable-periodic activation functions. Yet another worth noting work from this area is (Müller et al., 2022) with INR solution designed to effectively perform on modern computer architecture utilizing a simple data structure concept of hashmap to offer speed-oriented image representation with high fidelity of the outcomes.
Alternative approaches to INRs were presented in GaussianImage (Zhang et al., 2024). Instead of natural networks, the authors propose to approximate 2D images using 2D Gaussian components. In practice, GaussianImage is a 2D version of 3DGS (Kerbl et al., 2023) that uses 2D Gaussians instead of their 3D version and a simplified rendering procedure. Thanks to such a modification, the GaussianImage is invariant to the order of Gaussian components. Therefore, such a model is numerically efficient.
Figure 6: Flat 3D Gaussians can be utilized for 2D images in four distinct scenarios. Modifications demonstrate how the perspective of viewing a 2D image can change in each case. We establish a baseline using the GaMeS framework with Gaussian initialization on the XZ plane. Our results show that controlling Gaussian positions during learning enables various applications. The model employs a mirror setup during training, with the Amorphous configuration achieving the best results for image reconstruction and 3D analysis. The 2D model represents images on a single plane, allowing 2D physics engines like Taichi_elements to be used, but it does not support 3D modifications. In contrast, the Graphite model operates across multiple planes, making it ideal for 3D spatial reasoning and image combination.
@@
The representation and editing of objects using Gaussians is a well-explored topic in 3D graphics. In this field, meshes can be modified to simulate Gaussian editing (Guédon & Lepetit, 2024; Huang et al., 2024), or Gaussians can be directly parameterized and manipulated to achieve specific outcomes (Waczyńska et al., 2024b; a). This approach enables flexible and continuous deformations, offering an intuitive method for controlling object shapes and rendering properties, which has proven particularly useful in tasks like texture mapping, surface smoothing, and dynamic simulations.
@@ 3 MiraGe: Editable 2D Images using Gaussian Splatting
3D Gaussian Splatting
3DGS models 3D scene by a set of Gaussian components with color and opacity:
@@
During the rasterization stage, the 3DGS produces a sorted Gaussian list based on the projected depth information. Then, the $\alpha$-blending method is used to create the image. We refine the Gaussian parameters, color, and opacity in the training phase according to the reconstruction cost function. Training the Gaussian component is challenging. Consequently, the authors implement additional strategies for reducing and multiplying components. Gaussians with low opacity are removed, while those that change rapidly during optimization are multiplied. These strategies make the 3D Gaussian approach very efficient and capable of generating high-quality renders.
Figure 8: We integrate MiraGe with Material Point Methods (MPM) to achieve realistic alterations of 2D images. The initial column presents the original image, the subsequent two columns display renders captured midway through the simulation, and the final column shows the outcome of the full simulation. The colored lines in the last column trace the paths of randomly chosen points from the simulation.
GaMeS Parametrization of Gaussian Component
In MiraGe, we use flat Gaussian components in 3D space. In such a model we use Gaussian components with a covariance matrix $\Sigma$, factored as:
$\Sigma=RSSR^{T},$ (1)
where $R$ is the rotation matrix, and $S$ is a diagonal matrix containing the scaling parameters. However, we force one of the scale parameters to be zero. Consequently, we obtain a collection of flat Gaussian:
@@
where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$. In such a case, we can use GaMeS (Waczyńska et al., 2024b) parametrization to represent flat Gaussian by triangle-face mesh. This mapping is denoted by $\mathcal{T}(\cdot)$. When applied, this parametrization generates a set of triangles labeled as triangle soup.
To outline the GaMeS parameterization, consider a Gaussian component $\mathcal{N}({\bf m},R,S)$, characterized by the mean ${\bf m}$, the rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and the scaling matrix $S=\mathrm{diag}(\varepsilon,s_{2},s_{3})$. Then its face representation $\mathcal{N}(V)$ is based on a triangle:
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ (3)
with the vertices defined as: ${\bf v}_{1}={\bf m},\;\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\;\textrm{and}\;\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
$\mathcal{N}(\hat{\bf m},\hat{R},\hat{S})=\mathcal{N}(\mathcal{T}^{-1}(V))$ (4)
through the mean $\hat{\bf m}$, the rotation matrix $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$, and the scaling matrix $\hat{S}=\mathrm{diag}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3})$, where the parameters are defined by the following formulas:
$\hat{\bf m}={\bf v}_{1},\;\;\hat{\bf r}_{1}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$ (5)
$\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|},\;\;\hat{\bf r}_{3}=\mathrm{orth}({\bf v}_{3}-{\bf v}_{1};{\bf r}_{1},{\bf r}_{2}),$ (6)
$s_{1}=\varepsilon,\;\;\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|,\;\;\textrm{and}\;\;\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle.$ (7)
@@
The parametrization enables control over the Gaussians’ position, scale, and rotation by manipulating the underlying triangle mesh. Applying transformations to the triangle directly alters the corresponding Gaussian, as illustrated in Fig. 7.
MiraGe
In this work, we present an approach that leverages the concept of flat Gaussian distributions in 3D space to model a single 2D image as input. Our methodology is grounded in human visual perception. This perspective allows us to reframe the problem: instead of merely processing a pixel matrix, we interpret the images as objects with a fixed spatial configuration in a 3D environment.
@@
Figure 11: Comparison of PSNR obtained on a butterfly image from DIV2K dataset by MiraGe in comparison with GaussianImage (Zhang et al., 2024) and Gaussian Splatting (Kerbl et al., 2023). The lines with markers represent how the PSNR was changing during training. Different colors represent models trained with different numbers of Gaussians during initialization. The dashed black and orange lines represent the best results obtained during training by GaussianImage and GS, respectively. Vertical lines represent iteration, where MiraGe obtained better results than GaussianImage, the time in min:sec format above each line is the training time until this iteration. The experiment was performed on the rtx 4070 GPU.
We propose a method that situates the Gaussians within the $XZ$ plane, ensuring that the entire image remains visible under perspective projection. To achieve this, the possible range of $x$-values and $z$-values is calculated using the camera field of view. We first calculate the deviation from 00 on the $X$ axis using the similarity of triangles $\text{dev}_{z}=\text{cam}_{\text{dist}}\cdot\tan(0.5\cdot\text{Fov}_{\text{vert}})$, where $\text{cam}_{\text{dist}}$ and $\text{Fov}_{\text{vert}}$ are camera distance from the $XZ$ plane and camera field of view respectively. The deviation in the $X$ axis can be then computed by multiplying this value by the camera aspect ratio.
@@
Amorphous
The baseline approach how to control Gaussians is based on the classical GaMeS parametrization, initialized randomly on the $XZ$ plane, with the mean parameter’s $y$ coordinate set to zero:
@@
where ${\bf m}=[m_{1},0,m_{3}]$ $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$.
@@
Building on the promising outcomes of GaussianImage, we strategically anchored all Gaussians onto the $XZ$ plane. This configuration allows us to effectively translate the flat image geometry into a spatial framework, bridging perceptual intuition. We set the mean of these components to have zero in the second coordinate. Moreover, we use the projection of flat Gaussians on a 2D plane. Unfortunately, orthogonal projection can produce artifacts. Therefore, we use a rotation of Gaussian components to lay on the $XZ$ plane. Since we use flat Gaussians to extract such rotation we can use a rotation matrix between two vectors to align the vector in 3D (Markley, 1993). We use the notation $\mathrm{Rot}(a,b)$ for the rotation matrix.
MiraGe on 2D plane is define by set of 3G parameterized Gaussian components:
$\mathcal{G}=\{(\mathcal{N}([m_{1},0,m_{3}],[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]\mathrm{Rot}({\bf r}_{3},{\bf e}_{2}),\mathrm{diag}(\varepsilon,s_{2},s_{3})),\sigma_{i},c_{i})\},$ (9)
@@
where ${\bf e}_{2}=[0,1,0]$, ${\bf m}=[m_{1},0,m_{3}]$, $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$.
Graphite
@@
Unfortunately, 2D-MiraGe produces artifacts when we use modification in 3D space (see the third row in Fig. 6). Such an effect is coursed by the Gaussians, which appear randomly according to the camera position. To solve such a problem and obtain the possibility of 3D modifications, MiraGe allows the Gaussians to leave the $XZ$ plane:
$\mathcal{G}=\{(\mathcal{N}([m_{1},0,m_{3}]+\gamma{\bf e}_{2},[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]\mathrm{Rot}({\bf r}_{3},{\bf e}_{2}),\mathrm{diag}(\varepsilon,s_{2},s_{3})),\sigma_{i},c_{i})\},$ (10)
@@
where $\gamma$ is trainable parameter of translation scale along the vector ${\bf e}_{2}=[0,1,0]$, ${\bf m}=[m_{1},0,m_{3}]$, $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as: $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$. Such a model allows for the order of Gassians according to camera positions.
@@
By leveraging parameterized Gaussians, we achieved precise manipulation of 2D images directly within their native 2D space, enabling targeted edits of segmented regions and transformations of complete scenes in easier way. While this approach demonstrated substantial promise, we observed significant artifacts when extending manipulations into the 3D domain, particularly along the $Y$-axis.

### v2

@@ 1 Introduction
Figure 1: MiraGe encodes 2D images with parameterized Gaussians, enabling high-quality reconstruction and real-life-like modifications. The selected parts of the image can be transformed in 3D space, creating a 3D effect with a physics engine controlling movement and interactions.
@@
Gaussian Splatting (3DGS) framework models the structure of a 3D scene using Gaussian components (Kerbl et al., 2023). In the 2D domain, GaussianImage (Zhang et al., 2024) has shown promising results in image reconstruction by efficiently encoding images in the 2D space, with a strong focus on model efficiency and reduced training time. Unfortunately, GaussianImage does not support user-driven adjustments of scene objects, which is a key feature of 3DGS. While GaussianImage has explored image representation using 2D Gaussians primarily for data compression, our research highlights an additional benefit, i.e., the use of parameterized flat 3D Gaussians for editing 2D images. In our work, we address this by introducing the MiraGe model, which encodes 2D images through the lens of human perception, bridging the gap between 2D image representation and 3D spatial understanding (see Fig. 2).
Building on the foundational idea that humans intuitively can perform transformations on photographs–primarily through affine transformations and bending them beyond the 2D plane–we introduce a novel approach using flat Gaussians with GaMeS parametrization (Waczyńska et al., 2024b). This capability enables our model to support image editing in both 2D and 3D spaces. Notably, our framework simplifies often difficult perspective adjustments by allowing intuitive modifications directly within the third dimension (see Fig. 3).
Figure 3: Parameterized flat 3D Gaussians provide a powerful representation of 2D images, enabling flexible editing in 3D space. Triangle Soup can be animated using tools like Blender. The colored lines depict the motion paths of 10 randomly selected points during the simulation.
In addition to classical edits, our model has the unique capability of interfacing with physics engines, enabling applications that enhance the realism and immersiveness of animations (Jiang et al., 2024). MiraGe treats the physics engine as a black box and offers three distinct methods for controlling Gaussians, i.e., 2D, Amorphous and Graphite. For 2D representation (2D-MiraGe) we used Taichi_elements [^1] https://github.com/taichi-dev/taichi_elements, for 3D representation (Amorphous-MiraGe, Graphite-MiraGe) we use Blender [^2] https://www.blender.orgversion3.6. This flexibility makes our model highly applicable to various fields, such as computer graphics for populating spatial interfaces, where realistic, physics-factual 2D animations can be incorporated (Tadeja et al., 2023).
@@
Since high-quality image reconstruction is critical in animation, we compared MiraGe with other models, in particular with GaussianImage (see Fig. 5), showing our state-of-the-art performance in the image reconstruction task.
Figure 4: Two images were encoded using the MiraGe model on distinct planes within a 3D space. This setup allows for seamless integration of the encoded images, resulting in a collage-like composition. Moreover, the model facilitates editing capabilities, as illustrated here, with modifications to the background image (the rear plane).
It is worth highlighting that flat 3D Gaussians can be utilized for 2D images in four distinct scenarios, with modifications that emphasize how controlling the Gaussians during training affects the perspective of viewing each image (see Fig. 6). The following constitutes a list of our key contributions:
  • We introduce the MiraGe model, which represents 2D images using flat 3D Gaussian components, achieving state-of-the-art reconstruction quality.
@@
  • We integrate MiraGe with a physics engine, enabling physics-based modifications and interactions for both 2D and 3D environments.
Figure 5: Visual comparison of two Gaussian-based methods for 2D image reconstruction. From left to right, the columns display the ground truth image, the GaussianImage reconstruction, and the MiraGe reconstruction. The bottom row illustrates the differences between the ground truth image and the results of each method.
@@ 2 Related Works
One rapidly growing area in image reconstruction is Implicit Neural Representations (INRs), which have attracted significant attention for their ability to model continuous signals, such as images, through neural networks (Klocek et al., 2019). INRs encode spatial coordinates and map them to corresponding values, such as RGB color, allowing for highly compact and efficient representations (Xie et al., 2022). This has led to the development of several specialized models for image INR, such as SIREN (Sitzmann et al., 2020a), Fourier feature mapping (Tancik et al., 2020), and WIRE (Saragadam et al., 2023a). The growing field of research resulted in further improvements of already existing solutions, e.g. in (Liu et al., 2024), certain limitations of SIREN, namely the arising capacity-convergence gap, were successfully alleviated with the idea of variable-periodic activation functions. Yet another worth noting work from this area is (Müller et al., 2022) with INR solution designed to effectively perform on modern computer architecture utilizing a simple data structure concept of hashmap to offer speed-oriented image representation with high fidelity of the outcomes.
Figure 6: We demonstrate three approaches for Gaussian control: Amorphous, 2D, and Graphite. As a baseline, we utilize a single camera from the Amorphous setup. After applying perspective editing in 3D, the image shows noticeable deformation. In contrast, no deformation is observed when employing an additional camera. The model employs a mirror setup during training, with the Amorphous configuration achieving the best results for image reconstruction and 3D analysis. The 2D model represents images on a single plane. The Graphite model operates across multiple planes, making it ideal for 3D spatial reasoning and image combination.
Alternative approaches to INRs were presented in GaussianImage (Zhang et al., 2024). Instead of neural networks, the authors propose to approximate 2D images using 2D Gaussian components. In practice, GaussianImage is a 2D version of 3DGS (Kerbl et al., 2023) that uses 2D Gaussians instead of their 3D version and a simplified rendering procedure. Thanks to such a modification, the GaussianImage is invariant to the order of Gaussian components. Therefore, such a model is numerically efficient.
Figure 7: Integration of MiraGe with MPM enables realistic 2D image alterations. The first column shows the original image, the next two capture mid-simulation renders, and the last presents the final result with colored lines tracking point trajectories.
@@
The interactive image editing of 2D images has been widely explored in computer graphics. Here, some methods leverage the current advancements in generative models. For instance, (Pan et al., 2023) introduce DragGAN, enabling point-based manipulation of images by performing them on the underlying manifold of GAN, achieving realistic edits. Similarly, (Shi et al., 2023) propose DragDiffusion, which extends the previous framework to diffusion models, enhancing the control and applicability of image editing. On the other hand, (Jacobson et al., 2011) propose bounded biharmonic weights for linear blending, which produce smooth and intuitive deformation for handles of arbitrary topology. (Wang et al., 2015) further advances this field by proposing linear subspace design, unifying linear blend skinning and generalized barycentric coordinates to provide a practical way of controlling deformations.
The representation and editing of objects using Gaussians is a well-explored topic in 3D graphics. In this field, meshes can be modified to simulate Gaussian editing (Guédon & Lepetit, 2024; Huang et al., 2024), or Gaussians can be directly parameterized and manipulated to achieve specific outcomes (Waczyńska et al., 2024b, a). This approach enables flexible and continuous deformations, offering an intuitive method for controlling object shapes and rendering properties, which has proven particularly useful in tasks like texture mapping, surface smoothing, and dynamic simulations.
@@ 3 MiraGe: Editable 2D Images using Gaussian Splatting
Figure 8: Visual comparison of image editing techniques, demonstrating the effectiveness of representing 2D images with parameterized Gaussians applied to Triangle Soup. This approach enables highly realistic animations, achieving results comparable to those of generative models. Specifically, local editing operations preserve fine details, such as a dimple on a face, without affecting unrelated regions. Moreover, we can achieve precise manipulations, including subtle edits like closing a lion’s mouth, underscoring the flexibility and control inherent in our method.
@@
3D Gaussian Splatting 3DGS models 3D scene by a set of Gaussian components with color and opacity:
@@
Figure 9: Comparison of PSNR obtained on a butterfly image from DIV2K dataset by MiraGe, GaussianImage and GS. Vertical lines represent iteration, where MiraGe obtained better results than GaussianImage, the time in min:sec format above each line is the training time until this iteration.
During the rasterization stage, the 3DGS produces a sorted Gaussian list based on the projected depth information. Then, the $\alpha$-blending method is used to create the image. We refine the Gaussian parameters, color, and opacity in the training phase according to the reconstruction cost function. The optimal number of Gaussians required to represent a given object is not known a priori, and it is non-trivial to adjust the number of Gaussians. Hence, the initial number of Gaussians is a parameter of the method. The authors implement additional strategies for reducing and multiplying Gaussians. Gaussians with low opacity are removed, while those that change rapidly during optimization are multiplied. These strategies make the 3D Gaussian approach very efficient and capable of generating high-quality renders. We used this strategy to reconstruct 2D images, which distinguishes us from GaussianImage.
GaMeS Parametrization of Gaussian Component In MiraGe, we use flat Gaussian components in 3D space. In such a model we use Gaussian components with a covariance matrix $\Sigma$, factored as: $\Sigma=RSSR^{T}$, where $R$ is the rotation matrix, and $S$ is a diagonal matrix containing the scaling parameters. However, we force one of the scale parameters to be zero. Consequently, we obtain a collection of flat Gaussian:
@@
where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in R^{3}$. In such a case, we can use GaMeS parametrization to represent flat Gaussian by triangle-face mesh. This mapping is denoted by $\mathcal{T}(\cdot)$. When applied, this parametrization generates a set of triangles labeled as triangle soup.
To outline the GaMeS parametrization, consider a Gaussian component $\mathcal{N}({\bf m},R,S)$, characterized by the mean ${\bf m}$, the rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and the scaling matrix $S=\mathrm{diag}(\varepsilon,s_{2},s_{3})$. Then its face representation $\mathcal{N}(V)$ is based on a triangle: $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ with the vertices defined as: ${\bf v}_{1}={\bf m},\;\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\;\textrm{and}\;\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component $\mathcal{N}(\hat{\bf m},\hat{R},\hat{S})=\mathcal{N}(\mathcal{T}^{-1}(V))$ through the mean $\hat{\bf m}$, the rotation matrix $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$, and the scaling matrix $\hat{S}=\mathrm{diag}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3})$, where the parameters are defined by the following formulas:
$\hat{\bf m}={\bf v}_{1},\;\;\hat{\bf r}_{1}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$
$\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|},\;\;\hat{\bf r}_{3}=\mathrm{orth}({\bf v}_{3}-{\bf v}_{1};{\bf r}_{1},{\bf r}_{2}),$
$s_{1}=\varepsilon,\;\;\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|,\;\;\textrm{and}\;\;\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle.$
@@
The parametrization enables control over the Gaussians’ position, scale, and rotation by manipulating the underlying triangle mesh. Applying transformations to the triangle directly alters the corresponding Gaussian.
MiraGe In this work, we present an approach that leverages the concept of flat Gaussian distributions in 3D space to model a single 2D image as input. Our methodology is grounded in human visual perception. This perspective allows us to reframe the problem: instead of merely processing a pixel matrix, we interpret the images as objects with a fixed spatial configuration in a 3D environment.
@@
We propose a method that situates the Gaussians within the $XZ$ plane, ensuring that the entire image remains visible under perspective projection. To achieve this, the possible range of $x$-values and $z$-values is calculated using the camera field of view. We first calculate the deviation from 00 on the $X$ axis using the similarity of triangles $\text{dev}_{z}=\text{cam}_{\text{dist}}\cdot\tan(\frac{1}{2}\text{Fov}_{\text{vert}})$, where $\text{cam}_{\text{dist}}$ and $\text{Fov}_{\text{vert}}$ are camera distance from the $XZ$ plane and camera field of view respectively. The deviation in the $X$ axis can be then computed by multiplying this value by the camera aspect ratio.
@@
Amorphous The baseline approach how to control Gaussians is based on the classical GaMeS parametrization, initialized randomly on the $XZ$ plane, with the mean parameter’s $y$ coordinate set to zero:
@@
where ${\bf m}=[m_{1},0,m_{3}]$ $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$.
@@
Building on the promising outcomes of GaussianImage we anchored all Gaussians to the $XZ$ plane, translating flat image geometry into a spatial framework. We set the mean of these components to zero in the second coordinate. Moreover, we use the projection of flat Gaussians on a 2D plane. Unfortunately, orthogonal projection can produce artifacts. Therefore, we use a rotation of Gaussian components to lay on the $XZ$ plane. Since we use flat Gaussians to extract such rotation, we can use a rotation matrix between two vectors to align the vector in 3D (Markley, 1993). We use the notation $\mathrm{Rot}(a,b)$ for the rotation matrix.
MiraGe on 2D plane is define by set of 3D parameterized Gaussian components:
$\mathcal{G}=\{(\mathcal{N}({\bf m},R_{i}\,\mathrm{Rot}({\bf r}_{3},{\bf e}_{2}),S,\sigma_{i},c_{i})\},$
@@
where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, ${\bf e}_{2}=[0,1,0]$, ${\bf m}=[m_{1},0,m_{3}]$, and $R_{i}$ is the rotation matrix defined as $R_{i}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$.
@@
Graphite Unfortunately, 2D-MiraGe produces artifacts when we use modification in 3D space (see the third row in Fig. 6). Such an effect is coursed by the Gaussians, which appear randomly according to the camera position. To solve such a problem and obtain the possibility of 3D modifications, MiraGe allows the Gaussians to leave the $XZ$ plane:
$\mathcal{G}=\{(\mathcal{N}({\bf m}+\gamma{\bf e}_{2},R_{i}\,\mathrm{Rot}({\bf r}_{3},{\bf e}_{2}),S,\sigma_{i},c_{i})\},$
@@
where $\gamma$ is trainable parameter of translation scale along the vector ${\bf e}_{2}=[0,1,0]$, ${\bf m}=[m_{1},0,m_{3}]$, $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R_{i}$ is the rotation matrix defined as: $R_{i}=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$. Such a model allows for the order of Gassians according to camera positions.
@@
By leveraging parameterized Gaussians, we achieved precise manipulation of 2D images directly within their native 2D space, enabling targeted edits of segmented regions and transformations of complete scenes in an easier way. While this approach demonstrated substantial promise, we observed significant artifacts when extending manipulations into the 3D domain, particularly along the $Y$-axis, see first and last row in Fig 6.
Figure 10: We compare the animation capabilities of MiraGe a DragGAN model, highlighting the advantages of our Gaussian-based image representation. This approach enables highly realistic edits by not relying on generative techniques. Our method offers greater control during animation. For example, adjusting the position of a leg does not inadvertently alter facial features.
Table 1: Quantitative comparison with various baselines in PSNR and MS-SSIM. MiraGe gives state-of-the-art results. Model-$x$ denotes that the model was initialized with $x$ Gaussians.
Kodak dataset DIV2K dataset
PSNR $\uparrow$ MS-SSIM $\uparrow$ PSNR $\uparrow$ MS-SSIM $\uparrow$
WIRE 41.47 0.9939 35.64 0.9511
SIREN 40.83 0.9960 39.08 0.9958
I-NGP 43.88 0.9976 37.06 0.9950
NeuRBF 43.78 0.9964 38.60 0.9913
3DGS 43.69 0.9991 39.36 0.9979
GaussianImage-70k 44.08 0.9985 39.53 0.9975
GaussianImage-100k* 38.93 0.9948 41.48 0.9981
MiraGe-70k (our) 57.41 0.9998 53.22 0.9996
MiraGe-100k (our) 59.52 0.9999 54.54 0.9998